### v1

# Hierarchical Reasoning Model: A Critical Supplementary Material

Abstract
Transformers have demonstrated remarkable performance in natural language processing and related domains, as they largely focus on sequential, autoregressive next-token prediction tasks. Yet, they struggle in logical reasoning, not necessarily because of a fundamental limitation of these models, but possibly due to the lack of exploration of more creative uses, such as latent space and recurrent reasoning. An emerging exploration in this direction is the Hierarchical Reasoning Model Wang et al. (2025a), which introduces a novel type of recurrent reasoning in the latent space of transformers, achieving remarkable performance on a wide range of 2D reasoning tasks. Despite the promising results, this line of models is still at an early stage and calls for in-depth investigation. In this work, we perform a critical review on this class of models, examine key design choices and present intriguing variants that achieve significantly better performance on the Sudoku-Extreme and Maze-Hard tasks than previously reported. Our results also raise surprising observations and intriguing directions for further research.
1 Introduction
The Transformer architecture (Vaswani et al., 2017) has become the foundation of modern large language models (LLMs), powering systems such as BERT (Devlin et al., 2019), PaLM (Chowdhery et al., 2022) and GPT series (Brown et al., 2020; Achiam et al., 2023). Scaling these models has led to impressive progress across natural language understanding and generation tasks. Nevertheless, benchmarks specifically designed to probe reasoning — for example, GSM8K for arithmetic reasoning (Cobbe et al., 2021), BIG-Bench for diverse challenging tasks (Srivastava et al., 2022) and MATH for competition-level mathematics (Hendrycks et al., 2021) — have revealed limitations in systematic multi-step reasoning, even for the largest models.
To address these gaps, methods have been proposed to elicit reasoning through prompting or architectural modifications. Chain-of-thought prompting (Wei et al., 2022) and scratchpad supervision (Nye et al., 2021) train or prompt models to generate explicit intermediate steps, while program-aided approaches such as PAL (Gao et al., 2023) offload computation to external interpreters. At the same time, theoretical studies highlight structural limits of attention-only architectures: for example, Abbe et al. (2024) demonstrate a “locality barrier” that constrains global reasoning unless additional memory mechanisms are introduced. These insights echo findings from cognitive science, which emphasizes the hierarchical and compositional nature of human reasoning (Dedhe et al., 2023). Together, this body of work motivates the exploration of architectures and training paradigms that go beyond standard token-level generation, such as latent-space reasoning and recurrent refinement, which we examine in this paper, with an emphasis on the latest popular embodiment of these two ideas — Hierarchical Reasoning Model (HRM) (Wang et al., 2025a).
2 Latent Space Reasoning and Hierarchical Reasoning Model (HRM)
2.1 Latent Space vs. Token Space Reasoning
Token-space reasoning relies on generating explicit intermediate steps as natural-language tokens, as in chain-of-thought prompting (Wei et al., 2022). While such methods improve interpretability, they often introduce inefficiency: reasoning traces may be long, redundant, and brittle to errors in individual tokens (Nye et al., 2021; Kojima et al., 2022). Moreover, many benchmarks reveal that token-based reasoning, despite improvements, remains slow and sometimes unreliable on harder problems (Srivastava et al., 2022; Cobbe et al., 2021).
Latent-space reasoning takes a different approach, performing reasoning operations within the model’s hidden representations rather than explicit text. Recent work on latent chain-of-thought and related methods (Hao et al., 2024) shows that iterative refinement in latent space can yield more compact and flexible reasoning traces. These approaches reduce reliance on token generation, enabling richer intermediate states, faster inference, and in some cases improved accuracy (Hao et al., 2024). Latent-space reasoning thus provides a promising complement to explicit token-space methods.
2.2 HRM vs. Transformers
The Hierarchical Reasoning Model (HRM) (Wang et al., 2025a) is the latest popular model that performs latent-space reasoning. It introduces a dual-loop recurrent architecture inspired by cognitive science. It alternates between a fast low-level (L) module that executes candidate reasoning steps and a slower high-level (H) module that updates a more abstract plan. HRM iterates these cycles until a learned halting signal indicates completion, yielding effective depth far beyond its nominal layer count. Notably, HRM achieves strong performance on reasoning-heavy benchmarks such as Sudoku and ARC-AGI-1 (Chollet, 2019) without pretraining or explicit chain-of-thought supervision (Wang et al., 2025a). Despite the promising ideas and concepts introduced by the paper, it is unclear how much benefit these L and H modules bring, compared to a plain transformer. In our experiments in Figure 2(b), we show that even without H module, a plain 8-layer L-Module only HRM performs similarly to the original 4-layer L-module, 4-layer H-module HRM. This is consistent to what is reported by ARC-Foundation (2025), which says that the plain transformer performs similarly to HRM.
3 Recurrent Reasoning as a Type of Diffussion Model
3.1 HRM is a Diffusion Model
HRM attracted much attention upon release, as indicated by its 10k Github stars (Wang et al., 2025b) received within only two months. Its popularity can be attributed to its unique design and training procedure, as well as its strong performance. A key novelty is that HRM frameworks its recurrence in a non-Backpropagation Through Time (BPTT) setting and thus distinguishes itself sharply from any Recurrent Neural Networks (RNNs), giving people from the traditional language processing and RNN background a fresh impression.
However, HRM can be an interesting example model whose novelty actually varies by perspective: while it is quite innovative from a traditional RNN standpoint, it appears more routine from the alternative perspective we propose below:
HRM’s training algorithm — which replaces BPTT with a “one-step” gradient rule — turns out to be very similar to how diffusion models are trained. In a denoising diffusion process, one repeatedly applies a neural network to gradually denoise a sample, learning to map a noisy input back to the clean data (Ho et al., 2020; Rombach et al., 2022; Song et al., 2023). Similarly, HRM repeatedly refines an internal latent state towards the solution.
More formally, diffusion training uses example pairs to teach a network to recover the original data from a noisy version , where can be small (classical diffusion model (Ho et al., 2020; Rombach et al., 2022)) or large (consistency model (Song et al., 2023)). HRM’s one-step training can be viewed the same way: the model takes an intermediate reasoning state (analogous to ) and learns to map it directly to the final answer (analogous to ). The only substantive difference is the data: diffusion models add artificial noise to construct , whereas HRM’s latent state is an intermediate reasoning state generated by the model. Since HRM operates on latent space and learns directly the mapping from to , it corresponds precisely to “Latent Consistency Model” (LCM) (Song et al., 2024).
Consistency models (Song et al., 2023; 2024) can generate high-quality, visually appealing images with as few as 2-4 inference steps. As we will demonstrate later, a well-trained HRM on the Sudoku task also significantly reduces its inference steps, down to as few as 2 to 4 steps, which is surprising yet somewhat expected when viewing HRM as a LCM. From a task standpoint, few-step image generation appears plausible, as humans can all perform instant visual imagination. However, few-step reasoning on Sudoku is unexpected. It is intriguing to see how these models learn to solve these seemingly unrelated tasks in a similar manner when trained using the same framework.
3.2 BPTT vs. Diffusion: a Biological Perspective
A key motivation for HRM is biological plausibility: by avoiding BPTT, the authors claim a more brain-like learning rule. Once we recognize the parallel with diffusion models, it follows that any diffusion-like recurrent process could share this advantage. Diffusion models do not require backpropagation through multiple time steps; they train a feedforward mapping from one state to another independently. In HRM’s words, the one-step gradient has “constant memory” versus BPTT’s growing footprint. Diffusion training is inherently of this form: the model learns locally at each step without unfolding the entire recurrence in memory.
Therefore, diffusion models offer an equally biologically plausible paradigm for deep sequential computation. Generating a sample with a diffusion model is effectively a recurrent process (iteratively refining a latent) but it is trained with a feedforward loss. This suggests that mechanisms inspired by diffusion could provide a basis for “deep reasoning” in the brain without requiring implausible credit assignment. In summary, HRM’s replacement of BPTT is not unique: any latent consistency or diffusion model likewise avoids BPTT, pointing to a broader potential for biologically plausible recurrent learning. Further work should formalize this connection and explore how diffusion-like updates might operate in neural circuits.
However, it is important to note that avoiding BPTT does not equate to being fully biologically implementable. A deep feedforward model still requires credit assignment from later layers to earlier ones, where a few algorithms may be possible (e.g., Liao et al. (2024); Nøkland (2016); Lillicrap et al. (2016); Liao et al. (2015)). The biological plausibility of transformer blocks, on the other hand, remains an open question.
4 Adaptive Recurrent Reasoning for Both Training and Inference
4.1 Key Finding
We identify a critical mismatch in the original HRM implementation: while its Adaptive Computation Time (ACT) module learns to decide when to halt during training, evaluation ignores this policy and always executes the maximum number of reasoning steps. To address this, we modify HRM to use its learned halting policy during evaluation as well.
Our adjustment allows HRM at evaluation time to more faithfully reflect the model’s learned reasoning process. By accessing the learned ACT during inference, the model is able to decide earlier on that a puzzle is completed, preventing potential errors it could make if it were to complete its maximum number of reasoning steps. Indeed, as we observe in Figure 1 and Figure 2, our modified HRMs highly benefit from applying ACT halting during inference, greatly surpassing the original HRM performance. The similarity in performance between Figure 1, where we allow the 4-layer H- and L-module HRM to use ACT halting during evaluation, and Figure 2, where we allow an 8-layer L-module HRM to use ACT halting during evaluation, casts doubt on whether the biologically-inspired hierarchical H- and L- module design choice provides much benefit. In Figure 3, we observe that the model flexibly allocates computation steps across different puzzles, with some puzzles being solved in very few reasoning steps. Our findings also raise interesting questions about how models can reason, and how they differ from or align with human reasoning.
4.2 Methods
We build on the Hierarchical Reasoning Model (HRM) introduced by Wang et al. (2025a), which integrates the ACT module into a multi-step recurrent reasoning architecture. The HRM architecture consists of two hierarchical reasoning modules: an H module that processes global abstractions, updating less frequently, and an L module that handles local reasoning, which updates more frequently. The model also maintains an inner carry state that preserves computational context across reasoning iterations at every step. The model’s ACT module uses Q-learning to balance exploration (continuing reasoning) against exploitation (terminating with the current solution). At each step , a Q-head outputs values for halt and continue as . Halting occurs if (i) the maximum horizon is reached or (ii) once a stochastic minimum threshold has been satisfied.
Choosing halt gives reward , while continue yields
The loss at step combines prediction accuracy and stopping decisions:
In the released HRM source code (Wang et al., 2025b), however, while the ACT module is used during training, evaluation almost always executes for the full steps. This approach ignores the halting mechanism, ensuring uniform batch shapes but preventing evaluation from using the model’s learned stopping policy to its full potential. As a result, evaluation performance diverges from training performance, as the model under evaluation may take unnecessary extra steps even when a puzzle is already solved. By neglecting to use the ACT module to halt when the puzzle is already solved, potential computational efficiency gains and performance are compromised.
We modify HRM to apply halting learned during training to evaluation as well. Specifically, a sequence is completed when it either halts naturally or reaches the maximum step limit . We use rather than the stochastic comparison in the original HRM training implementation to determine whether to halt or continue the sequence, as evaluation does not benefit from using the exploratory logits.
Evaluation proceeds until all sequences in a batch have completed, whether through the ACT module halting the sequence or by evaluating the sequence up to the max step. Rather than running all the sequences up to the maximum step , as in the original HRM implementation, this modification allows evaluation to measure both task accuracy and the model’s ability to adaptively allocate computation.
Our changes ensure that evaluation tests the learned halting policy, not just maximum-depth reasoning of the model. We compare our HRM, where ACT is used in both training and evaluation, against the original HRM with fixed-step evaluation to measure accuracy improvements. Specifically, we train and evaluate this modified HRM, which utilizes its halting capabilities in the evaluation mode, on the released Sudoku-Extreme and Maze-Hard datasets.
4.3 Results
Remarkably, our modified HRM implementation, which uses ACT in both training and evaluation, largely outperforms the original HRM, achieving nearly perfect performance on the Sudoku-Extreme dataset compared to the original HRM’s performance (Figure 1). This shows that halting during evaluation is a crucial component for realizing the model’s reasoning potential. When the model continues reasoning after it has already produced a correct solution, additional steps can introduce errors. Over-reasoning can harm performance, as the model’s fixed-step evaluation can force it to keep making edits and predictions even when those updates are unnecessary. By contrast, allowing the ACT policy to halt during inference can help prevent this drift. We can observe more of the model’s flexible halting process, and its ability to reason in fewer steps during evaluation in Figure 3.
In Figure 2, we remove the H-module in the HRM and instead expand the L-module to 8 layers, while continuing to enable ACT during evaluation. This model, with the ability to use ACT during evaluation, also consistently achieves higher accuracy than the original HRM, while also requiring fewer steps on average to complete a puzzle. Despite this modified HRM using only a single 8-layer L-module, doing away with biologically-inspired hierarchical design choice of the fast-updating L-module and slow-updating H-module, our achieves similar performance as our experiments with the original L- and H-module HRM in Figure 1. From these similar performance results, it is unclear how much benefit the hierarchical L and H modules bring to the HRM.
4.4 Reasoning with Few Steps
Figure 3 provides concrete examples where the model halts after only a handful of reasoning steps, while still producing correct solutions. The model dynamically adjusts computation across puzzles and sometimes solves difficult puzzles with very short trajectories. This behavior contrasts with human reasoning, where solving a Sudoku puzzle typically involves many step-by-step constraint propagations until the grid is fully filled. Adaptive halting allows the model not only to avoid unnecessary computation but also to achieve correct solutions through compressed reasoning pathways, raising interesting questions about how neural models represent and execute reasoning.
5 Generalization to Other Datasets
5.1 Maze-Hard
To show that our approach generalizes to other datasets, we tested its performance on the Maze-Hard dataset. The results are shown in Figure 4. Our approach converges much faster and doubles the HRM’s performance within 30,000 training iterations.
5.2 ARC-AGI
ARC-AGI has a large board size and requires significantly more computational resources to run, which can take a few days on a single GPU. At the time of writing, we are actively running experiments on ARC-AGI. We will report results on ARC-AGI in a future revision of the paper.
6 Discussion
Our work highlights an interesting and strong phenomenon that as the model learns better, it solves the problem with fewer steps. It stands in stark contrast to the common belief that more reasoning steps during inference would lead to better performance. As the HRM paper repeatedly states in the “Inference-time scaling” section: “An effective neural model should exploit additional computational resources during inference to enhance performance … Models trained with a specific can generalize to higher computational limits during inference, leading to improved accuracy” (Wang et al., 2025a). As what we see in our results, if given the chance, a well trained model actually choose to solve problems with less steps, which is very counterintuitive and intriguing.
As another related investigation, ARC-Foundation (2025) published a blog post discussing many aspects about HRM, pointing out that the ACT plays an important factor in HRM’s good performance. However, ARC-Foundation (2025) does not observe much performance differences between any settings with ACT loop 4 or even 2. They also do not observe the importance of ACT during inference, nor any out-of-curve performances as we see. In fact, ARC-Foundation (2025) states: “This indicates that training with refinement loops is more important than using the refinement for inference.” This conclusion is exactly the opposite of what we show in this paper, where inference time ACT makes a significant difference.
7 Conclusion
In this paper, we perform a critical review on the recently popular reasoning model HRM. We show that HRM’s architecture can be largely simplified to a 8-layer plain transformer while maintaining good performance. We demystify HRM’s “one-step approximation of BPTT” by showing its equivalence to diffusion/consistency models. We explore variants of HRM and achieve drastically better performance than the original HRM model, raising the Sudoku-Extreme performance from previously reported 55% to above 90%. We highlight the interesting relationship between testing performance and testing reasoning steps, where the model learns to reason with fewer and fewer steps while achieving increasingly better performance. Visualization on Sudoku examples provides an intuitive demonstration of model’s surprising behavior, which differs greatly from human problem solving.
Acknowledgments
This work was supported by the K. Lisa Yang Integrative Computational Neuroscience (ICoN) Center, one of the six centers of the Yang Tan Collective at MIT. This work was also supported by the Center for Brains, Minds and Machines (CBMM), funded by NSF STC award CCF - 1231216.